Scenario 1 Consider the following game. Payoffs are in millions of dollars. Lawrence LLP Put Poison Pi Dump Cash Assets Buy Turbo Tech - Buy Zamboni Tech In Turbo Tech $100, -$1 $1.-$1 of Zamboni Tech $2, -$0.5 $0.5, -S0.5 ERS Corporation

1. What is the expected value for Player R playing R2 if they believe there is a 75 percent chance player C plays C1?

2.What is the expected value for Player R playing R1 if they believe there is a 50 percent chance player C plays C1?
Answer #1

Assuming Buy Turbo Tech is R1 strategy and Buy Zamboni Tech is R2 strategy for R(Row Player) & Put Poison Pill in turbo is C1 and Dump Cash Assets of Zamboni Tech is C2

Ans for 1)

Expeted value for Player R playing R2=0.75(1)+0.25(-0.5)=0.75-0.125=0.625 USD

Expected value for Player R playing R1=0.5(-100)+0.5(2)=-50+1=-49